OpenStudy (anonymous):

If x= 2 and r= 5, then y= ?

OpenStudy (lgbasallote):

\[x^2 + y^2 = r^2\] \[y^2 = r^2 - x^2\] \[y = \sqrt{r^2 - x^2}\] \[y = \sqrt{5^2 - 2^2}\]

OpenStudy (anonymous):

What about if r= 25 and y= 15 then x=?

OpenStudy (lgbasallote):

what do you think? :)

OpenStudy (anonymous):

Um, 25^2/15^2 ?

OpenStudy (anonymous):

Ohhh, nevermind. you totally told me up above! ^ lol. Thanks.

OpenStudy (lgbasallote):

haha yeah :)

OpenStudy (anonymous):

Sorry bout that, its that blonde coming through. :)

OpenStudy (lgbasallote):

lol i thought you were brunette